OpenStudy (anonymous):

what is 1/8 plus 8/9?

OpenStudy (bahrom7893):

\[\frac{1}{8}+\frac{8}9=\frac{9+64}{72}\]

OpenStudy (anonymous):

1/8 +8/9 get common denominators 9/72+64/72 73/72 or 1 1/72